I forgot how to do this:
3x squared = 2-1x

Found 2 solutions by Cromlix, solver91311:
Answer by Cromlix(4381)

3x^2 = 2 - x
3x^2 + x - 2 = 0
(3x - 2)(x + 1) = 0
3x - 2 = 0
3x = 2
x = 2/3
x + 1 = 0
x = -1
x = 2/3, -1
Hope this helps.
:-)

Answer by solver91311(24713)




Standard form:



Factor:



Apply the Zero Product Rule





Or





Check your work:





Checks.



Checks.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism